# spacing=nonfrench

### v1

@@ 1 Scaling by Thinking in Continuous Space
Early attempts at increasing the power of language models focused on scaling model size, a practice that requires extreme amounts of data and computation. More recently, researchers have explored ways to enhance the reasoning capability of models by scaling test time computation. The mainstream approach involves post-training on long chain-of-thought examples to develop the model's ability to verbalize intermediate calculations in its context window and thereby externalize thoughts.
Figure 1: We train a 3.5B parameter language model with depth recurrence. At test time, the model can iterate longer to use more compute and improve its performance. Instead of scaling test-time reasoning by ``verbalizing'' in long Chains-of-Thought, the model improves entirely by reasoning in latent space. Tasks that require less reasoning like OpenBookQA converge quicker than tasks like GSM8k, which effectively make use of more compute.
However, the constraint that expensive internal reasoning must always be projected down to a single verbalized next token appears wasteful; it is plausible that models could be more competent if they were able to natively ``think'' in their continuous latent space. One way to unlock this untapped dimension of additional compute involves adding a recurrent unit to a model. This unit runs in a loop, iteratively processing and updating its hidden state and enabling computations to be carried on indefinitely. While this is not currently the dominant paradigm, this idea is foundational to machine learning and has been (re-)discovered in every decade, for example as recurrent neural networks, diffusion models, and as universal or looped transformers.
@@ 2 Why Train Models with Recurrent Depth?
  • By constructing an architecture that is compute-heavy and small in parameter count, we hope to set a strong prior towards models that solve problems by ``thinking'', i.e. by learning meta-strategies, logic and abstraction, instead of memorizing. The strength of recurrent priors for learning complex algorithms has already been demonstrated in the ``deep thinking'' literature Schwarzschild et al. (2021b); Bansal et al. (2022); Schwarzschild et al. (2023).
@@
  • Section 7 visualizes what computation patterns emerge at scale with this architecture and training objective, showing that context-dependent behaviors emerge in latent space, such as ``orbiting'' when responding to prompts requiring numerical reasoning.
@@ 3.1 Macroscopic Design
This iterative architecture will look familiar to the other modern iterative modeling paradigm, diffusion models (Song & Ermon, 2019), especially latent diffusion models (Rombach et al., 2022). We ran several ablations with iterative schemes even more similar to diffusion models, such as $\mathbf{s}_{i}=R(\mathbf{e},\mathbf{s}_{i-1})+\mathbf{n}$ where $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma_{i}I_{n\cdot h})$, but find the injection of noise not to help in our preliminary experiments, which is possibly connected to our training objective. We also evaluated and $\mathbf{s}_{i}=R_{i}(\mathbf{e},\mathbf{s}_{i-1})$, i.e. a core block that takes the current step as input (Peebles & Xie, 2023), but find that this interacts badly with path independence, leading to models that cannot extrapolate.
@@ 3.2 Microscopic Design
Within each group, we broadly follow standard transformer layer design. Each block contains multiple layers, and each layer contains a standard, causal self-attention block using RoPE (Su et al., 2021) with a base of 500005000050000, and a gated SiLU MLP (Shazeer, 2020). We use RMSNorm (Zhang & Sennrich, 2019) as our normalization function. The model has learnable biases on queries and keys, and nowhere else. To stabilize the recurrence, we order all layers in the following ``sandwich'' format, using norm layers $n_{i}$, which is related, but not identical to similar strategies in (Ding et al., 2021; Team Gemma et al., 2024):
@@
While at small scales, most normalization strategies, e.g. pre-norm, post-norm and others, work almost equally well, we ablate these options and find that this normalization is required to train the recurrence at scale.
@@
In summary, we can summarize the architecture by the triplet $(l_{P},l_{R},l_{C})$, describing the number of layers in each stage, and by the number of recurrences $r$, which may vary in each forward pass. We train a number of small-scale models with shape (1,4,1)(1,4,1)( 1 , 4 , 1 ) and hidden size $h=1024$, in addition to a large model with shape (2,4,2)(2,4,2)( 2 , 4 , 2 ) and $h=5280$. This model has only 888 ``real'' layers, but when the recurrent block is iterated, e.g. 32 times, it unfolds to an effective depth of $2+4r+2=132$ layers, constructing computation chains that can be deeper than even the largest fixed-depth transformers (Levine et al., 2021; Merrill et al., 2022).
@@ 3.3 Training Objective
To keep computation and memory low at train time, we backpropagate through only the last $k$ iterations of the recurrent unit. This enables us to train with the heavy-tailed Poisson distribution $\Lambda$, as maximum activation memory and backward compute is now independent of $r$. We fix $k=8$ in our main experiments. At small scale, this works as well as sampling $k$ uniformly, but with set fixed, the overall memory usage in each step of training is equal. Note that the prelude block still receives gradient updates in every step, as its output $\mathbf{e}$ is injected in every step. This setup resembles truncated backpropagation through time, as commonly done with RNNs, although our setup is recurrent in depth rather than time (Williams & Peng, 1990; Mikolov et al., 2011).
@@ 4.1 Training Setup
Given access to only enough compute for a single large scale model run, we opted for a dataset mixture that maximized the potential for emergent reasoning behaviors, not necessarily for optimal benchmark performance. Our final mixture is heavily skewed towards code and mathematical reasoning data with (hopefully) just enough general webtext to allow the model to acquire standard language modeling abilities. All sources are publicly available. We provide an overview in Figure 4. Following Allen-Zhu & Li (2024), we directly mix relevant instruction data into the pretraining data. However, due to compute and time constraints, we were not able to ablate this mixture. We expect that a more careful data preparation could further improve the model's performance. We list all data sources in Appendix C.
@@
We construct a vocabulary of 655366553665536 tokens via BPE (Sennrich et al., 2016), using the implementation of Dagan (2024). In comparison to conventional tokenizer training, we construct our tokenizer directly on the instruction data split of our pretraining corpus, to maximize tokenization efficiency on the target domain. We also substantially modify the pre-tokenization regex (e.g. of Dagan et al. (2024)) to better support code, contractions and LaTeX. We include a <|begin_text|> token at the start of every document. After tokenizing our pretraining corpus, we pack our tokenized documents into sequences of length 4096. When packing, we discard document ends that would otherwise lack previous context, to fix an issue described as the ``grounding problem'' in Ding et al. (2024), aside from several long-document sources of mathematical content, which we preserve in their entirety.
@@
We scale the architecture described in Section 3, setting the layers to (2,4,2)(2,4,2)( 2 , 4 , 2 ), and train with a mean recurrence value of $\bar{r}=32$. We mainly scale by increasing the hidden size to $h=5280$, which yields 555555 heads of size of 969696. The MLP inner dimension is 179201792017920 and the RMSNorm $\varepsilon$ is $10^{-6}$. Overall this model shape has about 1.51.51.5B parameters in non-recurrent prelude and head, 1.51.51.5B parameters in the core recurrent block, and 0.50.50.5B in the tied input embedding.
@@ 4.2 Compute Setup and Hardware
We train this model using compute time allocated on the Oak Ridge National Lab's Frontier cluster, a large-scale HPE Cray system containing 8 x AMD MI250X GPU nodes. Nodes are connected via 4xHPE Slingshot NICs. The scheduling system is orchestrated through SLURM. We train in bfloat16 mixed precision using a PyTorch-based implementation (Zamirai et al., 2021).
Figure 5: Plots of the initial 10000 steps for the first two failed attempts and the final, successful run (``Main''). Note the hidden state collapse (middle) and collapse of the recurrence (right) in the first two failed runs, underlining the importance of our architecture and initialization in inducing a recurrent model and explain the underperformance of these runs in terms of pretraining loss (left).
@@
Nominally, each MI250X chip [^2] Technically, each node contains 4 dual-chip MI250X cards, but its main software stack (ROCm runtime) treats these chips as fully independent. achieves 192 TFLOP per GPU (AMD, 2021). For a single matrix multiplication, we measure a maximum achievable speed on these GPUs of 125 TFLOP/s on our software stack (ROCM 6.2.0, PyTorch 2.6 pre-release 11/02) (Bekman, 2023). Our implementation, using extensive PyTorch compilation and optimization of the hidden dimension to $h=5280$ achieves a single-node training speed of 108.75 TFLOP/s, i.e. 87% AFU (``Achievable Flop Utilization''). Due to the weight sharing inherent in our recurrent design, even our largest model is still small enough to be trained using only data (not tensor) parallelism, with only optimizer sharding (Rajbhandari et al., 2020) and gradient checkpointing on a per-iteration granularity. With a batch size of 1 per GPU we end up with a global batch size of 16M tokens per step, minimizing inter-GPU communication bandwidth.
@@
Training proceeded through 21 segments of up to 12 hours, which scheduled on Frontier mostly in early December 2024. We also ran a baseline comparison, where we train the same architecture but in a feedforward manner with only 1 pass through the core/recurrent block. This trained with the same setup for 180B tokens on 256 nodes with a batch size of 2 per GPU. Ultimately, we were able to schedule 795B tokens of pretraining of the main model. Due to our constant learning rate schedule, we were able to add additional segments ``on-demand'', when an allocation happened to be available.
@@
Table 1: Results on lm-eval-harness tasks zero-shot without chat template across various open-source models. We show ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021b), OpenBookQA (Mihaylov et al., 2018), PiQA (Bisk et al., 2020), SciQ (Johannes Welbl, 2017), and WinoGrande (Sakaguchi et al., 2021). We report normalized accuracy when provided.
@@
Ours, ($r=4$) 3.5B 0.8T 57.19 22.95 36.07 23.32 18.60 65.12 84.80 55.24
Ours, ($r=8$) 3.5B 0.8T 66.07 32.50 45.08 24.88 22.00 70.72 91. 5 55.64
Ours, ($r=16$) 3.5B 0.8T 68.43 34.38 48.65 29.21 24.00 73.99 93.60 57.77
@@ 4.3 Importance of Norms and Initializations at Scale
While the run obviously stopped improving in training loss (left plot), we find that this stall is due to the model's representation collapsing (Noci et al., 2022). The correlation of hidden states in the token dimension quickly goes to 1.0 (middle plot), meaning the model predicts the same hidden state for every token in the sequence. We find that this is an initialization issue that arises due to the recurrence operation. Every iteration of the recurrence block increases token correlation, mixing the sequence until collapse.
We attempt to fix this by introducing the embedding scale factor, switching back to a conventional pre-normalization block, and switching to the learned adapter. Initially, these changes appear to remedy the issue. Even though token correlation shoots close to 1.0 at the start (``Bad Run 2'', green), the model recovers after the first 150 steps. However, we quickly find that this training run is not able to leverage test-time compute effectively (right plot), as validation perplexity is the same whether 1 or 32 recurrences are used. This initialization and norm setup has led to a local minimum as the model has learned early to ignore the incoming state $\mathbf{s}$, preventing further improvements.
@@
With our successful configuration, training continues smoothly for the next 750B tokens without notable interruptions or loss spikes. We plot training loss and perplexity at different recurrence steps in Figure 6. In our material, we refer to the final checkpoint of this run as our ``main model'', which we denote as Huginn-0125333/hu: gIn/, transl. “thought”, is a raven depicted in Norse mythology. Corvids are surprisingly intelligent for their size, and and of course, as birds, able to unfold their wings at test-time..
@@ 5.2 Math and Coding Benchmarks
We also evaluate the model on math and coding. For math, we evaluate GSM8k (Cobbe et al., 2021) (as zero-shot and in the 8-way CoT setup), MATH ((Hendrycks et al., 2021a) with the Minerva evaluation rules (Lewkowycz et al., 2022)) and MathQA (Amini et al., 2019). For coding, we check MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021). Here we find that our model significantly surpasses all models except the latest OLMo-2 model in mathematical reasoning, as measured on GSM8k and MATH. On coding benchmarks the model beats all other general-purpose open-source models, although it does not outperform dedicated code models, such as StarCoder2 (Lozhkov et al., 2024), trained for several trillion tokens. We also note that while further improvements in language modeling are slowing down, as expected at this training scale, both code and mathematical reasoning continue to improve steadily throughout training, see Figure 8.
Figure 7: Performance on GSM8K CoT (strict match and flexible match), HellaSwag (acc norm.), and HumanEval (pass@1). As we increase compute, the performance on these benchmarks increases. HellaSwag only needs 888 recurrences to achieve near peak performance while other benchmarks make use of more compute.
@@ 5.3 Where does recurrence help most?
We evaluate ARC-C performance as a function of recurrence and number of few-shot examples in the context in Figure 9. Interestingly, without few-shot examples to consider, the model saturates in compute around 8-12 iterations. However, when more context is given, the model can reason about more information in context, which it does, saturating around 20 iterations if 1 example is provided, and 32 iterations, if 25-50 examples are provided, mirroring generalization improvements shown for recurrence (Yang et al., 2024a; Fan et al., 2025). Similarly, we see that if we re-evaluate OBQA in Table 5, but do not run the benchmark in "closed-book" format and rather provide a relevant fact, our recurrent model improves significantly almost closing the gap to OLMo-2. Intuitively this makes sense, as the recurrent models has less capacity to memorize facts but more capacity to reason about its context.

### v2

@@ 1 Scaling by Thinking in Continuous Space
Early attempts at increasing the power of language models focused on scaling model size, a practice that requires extreme amounts of data and computation. More recently, researchers have explored ways to enhance the reasoning capability of models by scaling test time computation. The mainstream approach involves post-training on long chain-of-thought examples to develop the model’s ability to verbalize intermediate calculations in its context window and thereby externalize thoughts.
Figure 1: We train a 3.5B parameter language model with depth recurrence. At test time, the model can iterate longer to use more compute and improve its performance. Instead of scaling test-time reasoning by “verbalizing” in long Chains-of-Thought, the model improves entirely by reasoning in latent space. Tasks that require less reasoning like OpenBookQA converge quicker than tasks like GSM8k, which effectively make use of more compute.
However, the constraint that expensive internal reasoning must always be projected down to a single verbalized next token appears wasteful; it is plausible that models could be more competent if they were able to natively “think” in their continuous latent space. One way to unlock this untapped dimension of additional compute involves adding a recurrent unit to a model. This unit runs in a loop, iteratively processing and updating its hidden state and enabling computations to be carried on indefinitely. While this is not currently the dominant paradigm, this idea is foundational to machine learning and has been (re-)discovered in every decade, for example as recurrent neural networks, diffusion models, and as universal or looped transformers.
@@ 2 Why Train Models with Recurrent Depth?
  • By constructing an architecture that is compute-heavy and small in parameter count, we hope to set a strong prior towards models that solve problems by “thinking”, i.e. by learning meta-strategies, logic and abstraction, instead of memorizing. The strength of recurrent priors for learning complex algorithms has already been demonstrated in the “deep thinking” literature Schwarzschild et al. (2021b); Bansal et al. (2022); Schwarzschild et al. (2023).
@@
  • Section 7 visualizes what computation patterns emerge at scale with this architecture and training objective, showing that context-dependent behaviors emerge in latent space, such as “orbiting” when responding to prompts requiring numerical reasoning.
@@ 3.1 Macroscopic Design
This iterative architecture will look familiar to the other modern iterative modeling paradigm, diffusion models (Song and Ermon, 2019), especially latent diffusion models (Rombach et al., 2022). We ran several ablations with iterative schemes even more similar to diffusion models, such as $\mathbf{s}_{i}=R(\mathbf{e},\mathbf{s}_{i-1})+\mathbf{n}$ where $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\sigma_{i}I_{n\cdot h})$, but find the injection of noise not to help in our preliminary experiments, which is possibly connected to our training objective. We also evaluated and $\mathbf{s}_{i}=R_{i}(\mathbf{e},\mathbf{s}_{i-1})$, i.e. a core block that takes the current step as input (Peebles and Xie, 2023), but find that this interacts badly with path independence, leading to models that cannot extrapolate.
@@ 3.2 Microscopic Design
Within each group, we broadly follow standard transformer layer design. Each block contains multiple layers, and each layer contains a standard, causal self-attention block using RoPE (Su et al., 2021) with a base of $50000$, and a gated SiLU MLP (Shazeer, 2020). We use RMSNorm (Zhang and Sennrich, 2019) as our normalization function. The model has learnable biases on queries and keys, and nowhere else. To stabilize the recurrence, we order all layers in the following “sandwich” format, using norm layers $n_{i}$, which is related, but not identical to similar strategies in (Ding et al., 2021; Team Gemma et al., 2024):
@@
While at small scales, most normalization strategies, e.g. pre-norm, post-norm and others, work almost equally well, we ablate these options and find that this normalization is required to train the recurrence at scale [^2] Note also that technically $n_{3}$ is superfluous, but we report here the exact norm setup with which we trained the final model..
@@
In summary, we can summarize the architecture by the triplet $(l_{P},l_{R},l_{C})$, describing the number of layers in each stage, and by the number of recurrences $r$, which may vary in each forward pass. We train a number of small-scale models with shape $(1,4,1)$ and hidden size $h=1024$, in addition to a large model with shape $(2,4,2)$ and $h=5280$. This model has only $8$ “real” layers, but when the recurrent block is iterated, e.g. 32 times, it unfolds to an effective depth of $1322+4r+2=$ 132 layers, constructing computation chains that can be deeper than even the largest fixed-depth transformers (Levine et al., 2021; Merrill et al., 2022).
@@ 3.3 Training Objective
To keep computation and memory low at train time, we backpropagate through only the last $k$ iterations of the recurrent unit. This enables us to train with the heavy-tailed Poisson distribution $\Lambda$, as maximum activation memory and backward compute is now independent of $r$. We fix $k=8$ in our main experiments. At small scale, this works as well as sampling $k$ uniformly, but with set fixed, the overall memory usage in each step of training is equal. Note that the prelude block still receives gradient updates in every step, as its output $\mathbf{e}$ is injected in every step. This setup resembles truncated backpropagation through time, as commonly done with RNNs, although our setup is recurrent in depth rather than time (Williams and Peng, 1990; Mikolov et al., 2011).
@@ 4.1 Training Setup
Given access to only enough compute for a single large scale model run, we opted for a dataset mixture that maximized the potential for emergent reasoning behaviors, not necessarily for optimal benchmark performance. Our final mixture is heavily skewed towards code and mathematical reasoning data with (hopefully) just enough general webtext to allow the model to acquire standard language modeling abilities. All sources are publicly available. We provide an overview in Figure 4. Following Allen-Zhu and Li (2024), we directly mix relevant instruction data into the pretraining data. However, due to compute and time constraints, we were not able to ablate this mixture. We expect that a more careful data preparation could further improve the model’s performance. We list all data sources in Appendix C.
@@
We construct a vocabulary of $65536$ tokens via BPE (Sennrich et al., 2016), using the implementation of Dagan (2024). In comparison to conventional tokenizer training, we construct our tokenizer directly on the instruction data split of our pretraining corpus, to maximize tokenization efficiency on the target domain. We also substantially modify the pre-tokenization regex (e.g. of Dagan et al. (2024)) to better support code, contractions and LaTeX. We include a <|begin_text|> token at the start of every document. After tokenizing our pretraining corpus, we pack our tokenized documents into sequences of length 4096. When packing, we discard document ends that would otherwise lack previous context, to fix an issue described as the “grounding problem” in Ding et al. (2024), aside from several long-document sources of mathematical content, which we preserve in their entirety.
@@
We scale the architecture described in Section 3, setting the layers to $(2,4,2)$, and train with a mean recurrence value of $\bar{r}=32$. We mainly scale by increasing the hidden size to $h=5280$, which yields $55$ heads of size of $96$. The MLP inner dimension is $17920$ and the RMSNorm $\varepsilon$ is $10^{-6}$. Overall this model shape has about $1.5$B parameters in non-recurrent prelude and head, $1.5$B parameters in the core recurrent block, and $0.5$B in the tied input embedding.
@@ 4.2 Compute Setup and Hardware
We train this model using compute time allocated on the Oak Ridge National Laboratory’s Frontier supercomputer. This HPE Cray system contains 9408 compute nodes with AMD MI250X GPUs, connected via 4xHPE Slingshot-11 NICs. The scheduling system is orchestrated through SLURM. We train in bfloat16 mixed precision using a PyTorch-based implementation (Zamirai et al., 2021).
Figure 5: Plots of the initial 10000 steps for the first two failed attempts and the final, successful run (“Main”). Note the hidden state collapse (middle) and collapse of the recurrence (right) in the first two failed runs, underlining the importance of our architecture and initialization in inducing a recurrent model and explain the underperformance of these runs in terms of pretraining loss (left).
@@
Nominally, each MI250X chip [^3] Technically, each node contains 4 dual-chip MI250X cards, but its main software stack (ROCm runtime) treats these chips as fully independent. achieves 192 TFLOP per GPU (AMD, 2021). For a single matrix multiplication, we measure a maximum achievable speed on these GPUs of 125 TFLOP/s on our software stack (ROCM 6.2.0, PyTorch 2.6 pre-release 11/02) (Bekman, 2023). Our implementation, using extensive PyTorch compilation and optimization of the hidden dimension to $h=5280$ achieves a single-node training speed of 108.75 TFLOP/s, i.e. 87% AFU (“Achievable Flop Utilization”). Due to the weight sharing inherent in our recurrent design, even our largest model is still small enough to be trained using only data (not tensor) parallelism, with only optimizer sharding (Rajbhandari et al., 2020) and gradient checkpointing on a per-iteration granularity. With a batch size of 1 per GPU we end up with a global batch size of 16M tokens per step, minimizing inter-GPU communication bandwidth.
@@
Training proceeded through 21 segments of up to 12 hours, which scheduled on Frontier mostly in early December 2024. We also ran a baseline comparison, where we train the same architecture but in a feedforward manner with only 1 pass through the core/recurrent block. This trained with the same setup for 180B tokens on 256 nodes with a batch size of 2 per GPU. Ultimately, we were able to schedule 795B tokens of pretraining of the main model. Due to our constant learning rate schedule, we were able to add additional segments “on-demand”, when an allocation happened to be available.
@@
Table 1: Results on lm-eval-harness tasks zero-shot across various open-source models. We show ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021a), OpenBookQA (Mihaylov et al., 2018), PiQA (Bisk et al., 2020), SciQ (Johannes Welbl, 2017), and WinoGrande (Sakaguchi et al., 2021). We report normalized accuracy when provided.
@@
Ours, ($r=4$) 3.5B 0.8T 49.07 27.99 43.46 23.39 28.20 64.96 80.00 55.24
Ours, ($r=8$) 3.5B 0.8T 65.11 35.15 58.54 25.29 35.40 73.45 92.10 55.64
Ours, ($r=16$) 3.5B 0.8T 69.49 37.71 64.67 31.25 37.60 75.79 93.90 57.77
@@ 4.3 Importance of Norms and Initializations at Scale
While the run obviously stopped improving in training loss (left plot), we find that this stall is due to the model’s representation collapsing (Noci et al., 2022). The correlation of hidden states in the token dimension quickly goes to 1.0 (middle plot), meaning the model predicts the same hidden state for every token in the sequence. We find that this is an initialization issue that arises due to the recurrence operation. Every iteration of the recurrence block increases token correlation, mixing the sequence until collapse.
We attempt to fix this by introducing the embedding scale factor, switching back to a conventional pre-normalization block, and switching to the learned adapter. Initially, these changes appear to remedy the issue. Even though token correlation shoots close to 1.0 at the start (“Bad Run 2”, green), the model recovers after the first 150 steps. However, we quickly find that this training run is not able to leverage test-time compute effectively (right plot), as validation perplexity is the same whether 1 or 32 recurrences are used. This initialization and norm setup has led to a local minimum as the model has learned early to ignore the incoming state $\mathbf{s}$, preventing further improvements.
@@
[…]
@@ 5.2 Math and Coding Benchmarks
We also evaluate the model on math and coding. For math, we evaluate GSM8k (Cobbe et al., 2021) (as zero-shot and in the 8-way CoT setup), MATH ((Hendrycks et al., 2021b) with the Minerva evaluation rules (Lewkowycz et al., 2022)) and MathQA (Amini et al., 2019). For coding, we check MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021). Here we find that our model significantly surpasses all models except the latest OLMo-2 model in mathematical reasoning, as measured on GSM8k and MATH. On coding benchmarks the model beats all other general-purpose open-source models, although it does not outperform dedicated code models, such as StarCoder2 (Lozhkov et al., 2024), trained for several trillion tokens. We also note that while further improvements in language modeling are slowing down, as expected at this training scale, both code and mathematical reasoning continue to improve steadily throughout training, see Figure 8.
Figure 7: Performance on GSM8K CoT (strict match and flexible match), HellaSwag (acc norm.), and HumanEval (pass@1). As we increase compute, the performance on these benchmarks increases. HellaSwag only needs $8$ recurrences to achieve near peak performance while other benchmarks make use of more compute.
@@ 5.3 Where does recurrence help most?
We evaluate ARC-C performance as a function of recurrence and number of few-shot examples in the context in Figure 9. Interestingly, without few-shot examples to consider, the model saturates in compute around 8-12 iterations. However, when more context is given, the model can reason about more information in context, which it does, saturating around 20 iterations if 1 example is provided, and 32 iterations, if 25-50 examples are provided, mirroring generalization improvements shown for recurrence (Yang et al., 2024a; Fan et al., 2025). Similarly, we see that if we re-evaluate OBQA in Table 5, but do not run the benchmark in the default lm-eval ”closed-book” format and rather provide a relevant fact, our recurrent model improves significantly almost closing the gap to OLMo-2. Intuitively this makes sense, as the recurrent models has less capacity to memorize facts but more capacity to reason about its context.